# Tradeoffs of Diagonal Fisher Information Matrix Estimators

### v2

@@ 7 Conclusion
We have analyzed two different estimators ^$\hat{\mathcal{I}}_{1}(\bm{\theta})$ and ^$\hat{\mathcal{I}}_{2}(\bm{\theta})$ for the diagonal entries of the FIM. The variances of these estimators are determined by both the non-linearly of the neural network and the moments of the exponential family. We find distinct scenarios where each estimator is preferable through empirically examining the estimator variances and bounds. We connect the investigated estimators with the empirical Fisher. Future directions include to extend the analysis to the FIM block diagonals (e.g. Martens & Grosse (2015); Ren & Goldfarb (2021)) and to adapt the current NN optimizers (e.g. Kingma & Ba (2015)) to consider the FIM variances.
8 Acknowledgements
The authors thank Frank Nielsen for their insightful feedback.
@@ 8 Acknowledgements
Appendix A: The Conditional Variances in Closed Form Pg A
$\hookrightarrow$ Appendix B: Off-Diagonal Variance Pg B
$\hookrightarrow$ Appendix C: Bounding the Trace Variance by Full Spectrum Pg C
$\hookrightarrow$ Appendix D: Second Central Moment of Categorical Distribution Pg D
$\hookrightarrow$ Appendix E: Empirical Results Continued Pg E
$\hookrightarrow$ Appendix F: “Empirical Fisher” Continued Pg F
@@
Appendix G: Derivation of Eqs. 5 and 6 Pg G
$\hookrightarrow$ Appendix H: Proof of Eq. 10 Pg H
@@
$\hookrightarrow$ Appendix K: Proof of Corollary 4.2 Pg K
$\hookrightarrow$ Appendix L: Proof of Proposition 4.3 Pg L
$\hookrightarrow$ Appendix M: Proof of Corollary 4.6 Pg M
$\hookrightarrow$ Appendix N: Proof of Theorem 4.7 Pg N
$\hookrightarrow$ Appendix O: Proof of Lemma 4.8 Pg O
$\hookrightarrow$ Appendix P: Proof of Proposition 5.1 Pg P
$\hookrightarrow$ Appendix Q: Proof of Theorem 5.2 Pg Q
$\hookrightarrow$ Appendix R: Proof of Lemma 6.1 Pg R
$\hookrightarrow$ Appendix S: Proof of Corollary F.1 Pg S
$\hookrightarrow$ Appendix T: Proof of Corollary F.2 Pg T
Appendix A The Conditional Variances in Closed Form
@@
Lemma A.1.
$\forall\bm{x}\in\Re^{I}$, $\forall{i}=1,\ldots,\dim(\bm{\theta})$,
$\mathcal{I}(\theta_{i}\,|\,\bm{x})=$ $\;\partial_{i}\bm{h}^{a}(\bm{x})\partial_{i}\bm{h}^{b}(\bm{x})\mathcal{I}_{ab}(\bm{h}\,|\,\bm{x}),$ (23)
@@
${\mathcal{V}}_{1}(\theta_{i}\,|\,\bm{x})=$ $\;\frac{1}{N}\partial_{i}\bm{h}^{a}(\bm{x})\partial_{i}\bm{h}^{b}(\bm{x})\partial_{i}\bm{h}^{c}(\bm{x})\partial_{i}\bm{h}^{d}(\bm{x})\cdot\left[\mathcal{K}^{p}_{abcd}(\bm{t}\,|\,\bm{x})-\mathcal{I}_{ab}(\bm{h}\,|\,\bm{x})\cdot\mathcal{I}_{cd}(\bm{h}\,|\,\bm{x})\right],$ (24)
${\mathcal{V}}_{2}(\theta_{i}\,|\,\bm{x})=$ $\;\frac{1}{N}\partial^{2}_{ii}{\bm{h}}^{a}(\bm{x})\partial^{2}_{ii}{\bm{h}}^{b}(\bm{x})\mathcal{I}_{ab}(\bm{h}\,|\,\bm{x}),$ (25)
where $\partial_{i}\bm{h}(\bm{x})\vcentcolon=\partial\bm{h}(\bm{x})/\partial\theta_{i}$; $\partial_{ij}\bm{h}(\bm{x})\vcentcolon=\partial^{2}\bm{h}(\bm{x})/\partial\theta_{i}\partial\theta_{j}$;
Proof.
The proof directly follows from Soen & Sun (2021, Equation 6), Soen & Sun (2021, Theorem 4), and Soen & Sun (2021, Theorem 6)
@@
Lemma A.1 shows that, for the former, ${\mathcal{V}}_{1}(\theta_{i}\,|\,\bm{x})$ only depends on 1st order derivatives; while ${\mathcal{V}}_{2}(\theta_{i}\,|\,\bm{x})$ only depends on the 2nd order derivatives. For the latter, ${\mathcal{V}}_{1}(\theta_{i}\,|\,\bm{x})$ depends on both the 2nd and 4th central moments of $\bm{t}(\bm{y})$; while ${\mathcal{V}}_{2}(\theta_{i}\,|\,\bm{x})$ only depends on the 2nd central moments.
@@
Given $\mathcal{I}_{ab}(\bm{h}\,|\,\bm{x})$ and $\partial_{i}\bm{h}^{a}(\bm{x})$, the computational complexity of all diagonal entries $\mathcal{I}(\theta_{i}\,|\,\bm{x})$ is $\mathcal{O}(T^{2}\dim(\bm{\theta}))$. If $\mathcal{K}^{p}_{abcd}(\bm{t}\,|\,\bm{x})$ and $\partial^{2}_{ii}{\bm{h}}^{a}(\bm{x})$ are given, then the computational complexity of the variances in Eqs. 24 and 25 is respectively $\mathcal{O}(T^{4}\dim(\bm{\theta}))$ and $\mathcal{O}(T^{2}\dim(\bm{\theta}))$. Each requires to evaluate a $T\times\dim(\bm{\theta})$ matrix, either $\partial_{i}\bm{h}^{a}(\bm{x})$ or $\partial^{2}_{ii}{\bm{h}}^{a}(\bm{x})$ — which can be expensive to calculate for the latter. This is why we need efficient estimators and / or bounds for the tensors on the LHS of Eqs. 23, 24 and 25.
Appendix B Off-Diagonal Variance

### v3

@@ 7 Conclusion
We have analyzed two different estimators ^$\hat{\mathcal{I}}_{1}(\bm{\theta})$ and ^$\hat{\mathcal{I}}_{2}(\bm{\theta})$ for the diagonal entries of the FIM. The variances of these estimators are determined by both the non-linearly of the neural network and the moments of the exponential family. We have identified distinct scenarios on which estimator is preferable. For example, ReLU networks can only apply ^$\hat{\mathcal{I}}_{1}(\bm{\theta})$ due to a lack of smoothness. As another example, ^$\hat{\mathcal{I}}_{2}(\bm{\theta})$ has zero variance in the last layer and thus is always preferable than ^$\hat{\mathcal{I}}_{1}(\bm{\theta})$. Similarly, in the second last layer, ^$\hat{\mathcal{I}}_{2}(\bm{\theta})$ has a simple closed form and potentially preferable for neurons in their linear regions (see Remark 4.5). In general, one has to apply Theorem 4.1 based on their specific neural network and settings and choose the estimator with the smaller variance. Our results suggest that, from a variance perspective, uniformly utilizing one of the FIM estimators ^$\hat{\mathcal{I}}_{j}(\bm{\theta})$ is often suboptimal in NNs. Our work has further extended from analyzing the conditional FIM estimators ^$\hat{\mathcal{I}}_{j}(\bm{\theta}\,|\,\bm{x})$ to the joint FIM estimators ^$\hat{\mathcal{I}}_{j}(\bm{\theta})$; and we have examined the relationship between the investigated estimators and the empirical Fisher. Future directions include extending the analysis of the variance of FIM estimators to block diagonals (e.g. Martens and Grosse (2015); Ren and Goldfarb (2021)) and adapting current NN optimizers (e.g. Kingma and Ba (2015)) to incorporate the variance of FIM estimators.
Acknowledgements
The authors thank Frank Nielsen, James C. Spall, and the anonymous reviewers for their insightful feedback and many constructive comments.
@@ Acknowledgements
Appendix A: Natural Gradient Toy Data Example Pg A
$\hookrightarrow$ Appendix B: The Conditional Variances in Closed Form Pg B
$\hookrightarrow$ Appendix C: Off-Diagonal Variance Pg C
$\hookrightarrow$ Appendix D: Bounding the Trace Variance by Full Spectrum Pg D
$\hookrightarrow$ Appendix E: Second Central Moment of Categorical Distribution Pg E
$\hookrightarrow$ Appendix F: Empirical Results Continued Pg F
$\hookrightarrow$ Appendix G: “Empirical Fisher” Continued Pg G
@@
Appendix H: Derivation of Eq. 3 Using Log-Partition Function Derivatives Pg H
@@
$\hookrightarrow$ Appendix K: Proof of Eq. 9 Pg K
$\hookrightarrow$ Appendix L: Proof of Corollary 4.2 Pg L
$\hookrightarrow$ Appendix M: Proof of Proposition 4.3 Pg M
$\hookrightarrow$ Appendix N: Proof of Corollary 4.6 Pg N
$\hookrightarrow$ Appendix O: Proof of Theorem 4.7 Pg O
$\hookrightarrow$ Appendix P: Proof of Lemma 4.8 Pg P
$\hookrightarrow$ Appendix Q: Proof of Proposition 5.1 Pg Q
$\hookrightarrow$ Appendix R: Proof of Theorem 5.2 Pg R
$\hookrightarrow$ Appendix S: Proof of Lemma 6.1 Pg S
$\hookrightarrow$ Appendix T: Proof of Corollary G.1 Pg T
$\hookrightarrow$ Appendix U: Proof of Corollary G.2 Pg U
Appendix A Natural Gradient Toy Data Example
The following section describes the data and models of Fig. 1. In general, the toy data and models constructed consists of taking 1D output setting presented by Section 5, where the NN $h_{\bm{\theta}}(\bm{x})$ is a linear function.
A.I Data
The 2D input data $\bm{x}\in\Re^{2}$ is sampled from a simple isotropic centered Gaussian $\mathcal{N}(\bm{0},\bm{I})$. A linear response variable $a\in\Re$ is defined by the following:
$a=\bm{\theta}_{\rm true}^{\top}\bm{x};\quad\textrm{where $\bm{\theta}_{\rm true}=(1,1)$.}$
The outputs of $y$ for the cases of regression and classification are differentiated by how $a$ is used in sampling:
$y_{\textrm{regression}}$ $\sim\mathcal{N}(\mu=1,\sigma=1)$
$y_{\textrm{classification}}$ $\sim\mathrm{Bern}(p=\sigma(a)),$
where $\sigma(z)=(1+\exp(-z))$ is the logistic function.
A.II Model
The model $h_{\bm{\theta}}(\bm{x})=\bm{\theta}^{\top}\bm{x}$ consists of a linear function; and the exponential family Eq. 1 is chosen to be a 1D isotropic Gaussian and binary multinomial distribution (Bernoulli) for regression and classification, respectively. This corresponds to Section 5 for 1D outputs. Notice that the model exactly matches the data generating function.
A.III Training
Natural gradient descent (NGD) is taken using both ^$\hat{\mathcal{I}}_{1}(\bm{\theta})$ and ^$\hat{\mathcal{I}}_{2}(\bm{\theta})$. The estimated FIM utilize only a single $y\,|\,\bm{x}$ sample for each input $\bm{x}$. We use a learning rate of $\eta=0.01$ over $256$ epochs. A training set of $256$ data points are sampled. At each iteration of NGD, we sample $4$ random points from the training set for the update. The test loss is evaluated on a test set of $4096$ data points sampled.
A.IV Variance Plot of Example
Larger version of Fig. 1 with additional variance sum plotted over time is given by Fig. I. Note that variance sum is including off diagonals. Further note that the variance is calculated over joint sample in $(\bm{x},y)$.
Figure I: Extended version of Fig. 1 with the sum of variance of FIM estimators over epochs.
A.V Other Seeds
We further present other random seed of the teaser plot in Figs. II, III and IV.
Figure II: Fig. I over different randomizations (a).
Figure III: Fig. I over different randomizations (b).
Figure IV: Fig. I over different randomizations (c).
Appendix B The Conditional Variances in Closed Form
@@
Proof of Lemma 3.1.
The proof directly follows from Soen and Sun [2021, Equation 6], Soen and Sun [2021, Theorem 4], and Soen and Sun [2021, Theorem 6]. In what follows, we provide a proof of the Lemma utilizing the notation of this paper for completeness. We prove the statement one equation at a time.
For Eq. 4, we consider the following computation.
$\mathcal{I}(\theta_{i}\,|\,\bm{x})$
$=\mathop{\mathbb{E}}_{p(\bm{y}\,|\,\bm{x};\bm{\theta})}\left[\frac{\partial\log p(\bm{y}\,|\,\bm{x};\bm{\theta})}{\partial\bm{\theta}}\frac{\partial\log p(\bm{y}\,|\,\bm{x};\bm{\theta})}{\partial\bm{\theta}^{\top}}\right]$
$=\mathop{\mathbb{E}}_{p(\bm{y}\,|\,\bm{x};\bm{\theta})}\left[\frac{\partial\left(\bm{t}^{\top}(\bm{y})\bm{h}_{\bm{\theta}}(\bm{x})-F(\bm{h}_{\bm{\theta}}(\bm{x}))\right)}{\partial\bm{\theta}}\frac{\partial\left(\bm{t}^{\top}(\bm{y})\bm{h}_{\bm{\theta}}(\bm{x})-F(\bm{h}_{\bm{\theta}}(\bm{x}))\right)}{\partial\bm{\theta}^{\top}}\right]$
@@
$=\mathop{\mathbb{E}}_{p(\bm{y}\,|\,\bm{x};\bm{\theta})}\left[\left(\frac{\partial\bm{h}_{\bm{\theta}}(\bm{x})}{\partial\bm{\theta}}\right)^{\top}\left(\bm{t}(\bm{y})-\frac{\partial F(\bm{h})}{\partial\bm{h}}\bigg{|}_{{\bm{h}}=\bm{h}_{\bm{\theta}}(\bm{x})}\right)\left(\bm{t}(\bm{y})-\frac{\partial F(\bm{h})}{\partial\bm{h}}\bigg{|}_{{\bm{h}}=\bm{h}_{\bm{\theta}}(\bm{x})}\right)^{\top}\left(\frac{\partial\bm{h}_{\bm{\theta}}(\bm{x})}{\partial\bm{\theta}^{\top}}\right)\right]$
$=\mathop{\mathbb{E}}_{p(\bm{y}\,|\,\bm{x};\bm{\theta})}\left[\left(\frac{\partial\bm{h}_{\bm{\theta}}(\bm{x})}{\partial\bm{\theta}}\right)^{\top}\left(\bm{t}(\bm{y})-{\bm{\eta}}(\bm{x})\right)\left(\bm{t}(\bm{y})-{\bm{\eta}}(\bm{x})\right)^{\top}\left(\frac{\partial\bm{h}_{\bm{\theta}}(\bm{x})}{\partial\bm{\theta}^{\top}}\right)\right]$
$=\left(\frac{\partial\bm{h}_{\bm{\theta}}(\bm{x})}{\partial\bm{\theta}}\right)^{\top}\left(\mathop{\mathbb{E}}_{p(\bm{y}\,|\,\bm{x};\bm{\theta})}\left[\left(\bm{t}(\bm{y})-{\bm{\eta}}(\bm{x})\right)\left(\bm{t}(\bm{y})-{\bm{\eta}}(\bm{x})\right)^{\top}\right]\right)\left(\frac{\partial\bm{h}_{\bm{\theta}}(\bm{x})}{\partial\bm{\theta}^{\top}}\right)$
$=\left(\frac{\partial\bm{h}_{\bm{\theta}}(\bm{x})}{\partial\bm{\theta}}\right)^{\top}\mathcal{I}(\bm{h}\,|\,\bm{x})\left(\frac{\partial\bm{h}_{\bm{\theta}}(\bm{x})}{\partial\bm{\theta}^{\top}}\right).$
Using Einstein notation and restricting the partial derivative to a component of $\bm{\theta}$ yields the desired result.
For Eq. 5, we shorthand $\bm{\delta}(\bm{y})=\bm{t}(\bm{y})-\bm{\eta}(\bm{x})$. Note that the ^$\hat{\mathcal{I}}_{1}(\theta_{i}\,|\,\bm{x})$ estimator can be written as follows:
^$\hat{\mathcal{I}}_{1}(\theta_{i}\,|\,\bm{x})$ $=\frac{1}{N}\sum_{k=1}^{N}\left(\frac{\partial F(\bm{h}(\bm{x}))}{\partial\theta_{i}}-\frac{\partial\bm{h}^{a}(\bm{x})}{\partial\theta_{i}}\cdot\bm{t}_{a}(\bm{y}_{k})\right)^{2}$
$=\frac{1}{N}\sum_{k=1}^{N}\left(\frac{\partial\bm{h}^{a}(\bm{x})}{\partial\theta_{i}}\cdot\bm{\eta}_{a}(\bm{x})-\frac{\partial\bm{h}^{a}(\bm{x})}{\partial\theta_{i}}\cdot\bm{t}_{a}(\bm{y}_{k})\right)^{2}$
$=\frac{1}{N}\sum_{k=1}^{N}\partial\bm{h}^{a}(\bm{x})\partial\bm{h}^{b}(\bm{x}){\bm{\delta}}_{a}(\bm{y}_{k}){\bm{\delta}}_{b}(\bm{y}_{k}).$
Thus, we have
${\mathcal{V}}_{1}(\theta_{i}\,|\,\bm{x})$
$=\mathrm{Var}\left(\hat{\mathcal{I}}_{1}(\theta_{i}\,|\,\bm{x})\right)$
$=\mathrm{Var}\left(\frac{1}{N}\sum_{k=1}^{N}\partial\bm{h}^{a}(\bm{x})\partial\bm{h}^{b}(\bm{x}){\bm{\delta}}_{a}(\bm{y}_{k}){\bm{\delta}}_{b}(\bm{y}_{k})\right)$
$=\frac{1}{N}\cdot\mathrm{Var}\left(\partial\bm{h}^{a}(\bm{x})\partial\bm{h}^{b}(\bm{x}){\bm{\delta}}_{a}(\bm{y}_{k}){\bm{\delta}}_{b}(\bm{y}_{k})\right)$
$=\frac{1}{N}\cdot\left(\mathop{\mathbb{E}}_{p(\bm{y}\,|\,\bm{x};\bm{\theta})}\left[\left(\partial\bm{h}^{a}(\bm{x})\partial\bm{h}^{b}(\bm{x}){\bm{\delta}}_{a}(\bm{y}){\bm{\delta}}_{b}(\bm{y})\right)^{2}\right]-\mathop{\mathbb{E}}_{p(\bm{y}\,|\,\bm{x};\bm{\theta})}\left[\partial\bm{h}^{a}(\bm{x})\partial\bm{h}^{b}(\bm{x}){\bm{\delta}}_{a}(\bm{y}){\bm{\delta}}_{b}(\bm{y})\right]^{2}\right).$
Let us compute each of these terms.
$\mathop{\mathbb{E}}_{p(\bm{y}\,|\,\bm{x};\bm{\theta})}\left[\left(\partial\bm{h}^{a}(\bm{x})\partial\bm{h}^{b}(\bm{x}){\bm{\delta}}_{a}(\bm{y}){\bm{\delta}}_{b}(\bm{y})\right)^{2}\right]$
$=\mathop{\mathbb{E}}_{p(\bm{y}\,|\,\bm{x};\bm{\theta})}\left[\partial\bm{h}^{a}(\bm{x})\partial\bm{h}^{b}(\bm{x})\partial\bm{h}^{c}(\bm{x})\partial\bm{h}^{d}(\bm{x}){\bm{\delta}}_{a}(\bm{y}){\bm{\delta}}_{b}(\bm{y}){\bm{\delta}}_{c}(\bm{y}){\bm{\delta}}_{d}(\bm{y})\right]$
$=\partial\bm{h}^{a}(\bm{x})\partial\bm{h}^{b}(\bm{x})\partial\bm{h}^{c}(\bm{x})\partial\bm{h}^{d}(\bm{x})\mathop{\mathbb{E}}_{p(\bm{y}\,|\,\bm{x};\bm{\theta})}\left[{\bm{\delta}}_{a}(\bm{y}){\bm{\delta}}_{b}(\bm{y}){\bm{\delta}}_{c}(\bm{y}){\bm{\delta}}_{d}(\bm{y})\right]$
$=\partial\bm{h}^{a}(\bm{x})\partial\bm{h}^{b}(\bm{x})\partial\bm{h}^{c}(\bm{x})\partial\bm{h}^{d}(\bm{x})\mathcal{K}_{abcd}^{p}(\bm{t}\,|\,\bm{x}).$
And,
$\left(\mathop{\mathbb{E}}_{p(\bm{y}\,|\,\bm{x};\bm{\theta})}\left[\partial\bm{h}^{a}(\bm{x})\partial\bm{h}^{b}(\bm{x}){\bm{\delta}}_{a}(\bm{y}){\bm{\delta}}_{b}(\bm{y})\right]\right)^{2}$
$=\left(\partial\bm{h}^{a}(\bm{x})\partial\bm{h}^{b}(\bm{x})\mathop{\mathbb{E}}_{p(\bm{y}\,|\,\bm{x};\bm{\theta})}\left[{\bm{\delta}}_{a}(\bm{y}){\bm{\delta}}_{b}(\bm{y})\right]\right)^{2}$
$=\left(\partial\bm{h}^{a}(\bm{x})\partial\bm{h}^{b}(\bm{x})\mathcal{I}_{ab}(\bm{h}\,|\,\bm{x})\right)^{2}$
$=\partial\bm{h}^{a}(\bm{x})\partial\bm{h}^{b}(\bm{x})\partial\bm{h}^{c}(\bm{x})\partial\bm{h}^{d}(\bm{x})\mathcal{I}_{ab}(\bm{h}\,|\,\bm{x})\mathcal{I}_{cd}(\bm{h}\,|\,\bm{x})$
$=\partial\bm{h}^{a}(\bm{x})\partial\bm{h}^{b}(\bm{x})\partial\bm{h}^{c}(\bm{x})\partial\bm{h}^{d}(\bm{x})\left(\mathcal{I}(\bm{h}\,|\,\bm{x})\otimes\mathcal{I}(\bm{h}\,|\,\bm{x})\right)_{abcd}$
Simplifying all term yields the result as required.
Finally, for Eq. 6 we consider the following simplification of the estimator.
^$\hat{\mathcal{I}}_{2}(\theta_{i}\,|\,\bm{x})$ $=\frac{1}{N}\sum_{k=1}^{N}\left(\frac{\partial^{2}F(\bm{h}(\bm{x}))}{\partial^{2}\theta_{i}}-\frac{\partial^{2}\bm{h}^{a}(\bm{x})}{\partial^{2}\theta_{i}}\cdot\bm{t}_{a}(\bm{y}_{k})\right)$
$=\frac{1}{N}\sum_{k=1}^{N}\left(\frac{\partial}{\partial\theta_{i}}\left(\frac{\partial\bm{h}^{a}(\bm{x})}{\partial\theta_{i}}\cdot\bm{\eta}_{a}(\bm{x})\right)-\frac{\partial^{2}\bm{h}^{a}(\bm{x})}{\partial^{2}\theta_{i}}\cdot\bm{t}_{a}(\bm{y}_{k})\right)$
$=\frac{1}{N}\sum_{k=1}^{n}\left(\frac{\partial\bm{h}^{a}(\bm{x})}{\partial\theta_{i}}\cdot\frac{\partial\bm{\eta}_{a}(\bm{x})}{\partial\theta_{i}}+\frac{\partial^{2}\bm{h}^{a}(\bm{x})}{\partial^{2}\theta_{i}}\cdot\bm{\eta}_{a}(\bm{x})-\frac{\partial^{2}\bm{h}^{a}(\bm{x})}{\partial^{2}\theta_{i}}\cdot\bm{t}_{a}(\bm{y}_{k})\right)$
$=\frac{1}{N}\sum_{k=1}^{n}\left(\partial_{i}\bm{h}^{a}(\bm{x})\cdot\frac{\partial\bm{\eta}_{a}(\bm{x})}{\partial\theta_{i}}-\partial_{i}^{2}\bm{h}^{a}(\bm{x})\cdot\bm{\delta}_{a}(\bm{y}_{k})\right)$
$=\frac{1}{N}\sum_{k=1}^{n}\left(\partial_{i}\bm{h}^{a}(\bm{x})\cdot\partial_{i}\bm{h}^{b}(\bm{x})\cdot\mathcal{I}_{ab}(\bm{h}\,|\,\bm{x})-\partial_{i}^{2}\bm{h}^{a}(\bm{x})\cdot\bm{\delta}_{a}(\bm{y}_{k})\right)$
$=\partial_{i}\bm{h}^{a}(\bm{x})\cdot\partial_{i}\bm{h}^{b}(\bm{x})\cdot\mathcal{I}_{ab}(\bm{h}\,|\,\bm{x})-\frac{1}{N}\sum_{k=1}^{n}\left(\partial_{i}^{2}\bm{h}^{a}(\bm{x})\cdot\bm{\delta}_{a}(\bm{y}_{k})\right),$
where the last line follows from Soen and Sun [2021, Lemma 2] (a result of $p(\bm{y}\,|\,\bm{x};\bm{\theta})$ following an exponential family, see Amari [2016]).
Notice that the first quantity is a constant w.r.t. the randomness of $\bm{y}_{k}$. As such, we can simplify the variance calculation as follows.
${\mathcal{V}}_{2}(\theta_{i}\,|\,\bm{x})=\mathrm{Var}\left(\hat{\mathcal{I}}_{2}(\theta_{i}\,|\,\bm{x})\right)$
$=\mathrm{Var}\left(\partial_{i}\bm{h}^{a}(\bm{x})\cdot\partial_{i}\bm{h}^{b}(\bm{x})\cdot\mathcal{I}_{ab}(\bm{h}\,|\,\bm{x})-\frac{1}{N}\sum_{k=1}^{n}\left(\partial_{i}^{2}\bm{h}^{a}(\bm{x})\cdot\bm{\delta}_{a}(\bm{y}_{k})\right)\right)$
$=\mathrm{Var}\left(\frac{1}{N}\sum_{k=1}^{n}\left(\partial_{i}^{2}\bm{h}^{a}(\bm{x})\cdot\bm{\delta}_{a}(\bm{y}_{k})\right)\right)$
$=\frac{1}{N}\mathrm{Var}\left(\partial_{i}^{2}\bm{h}^{a}(\bm{x})\cdot\bm{\delta}_{a}(\bm{y})\right)$
$=\frac{1}{N}\cdot\left(\mathop{\mathbb{E}}_{p(\bm{y}\,|\,\bm{x};\bm{\theta})}\left[\left(\partial_{i}^{2}\bm{h}^{a}(\bm{x})\cdot\bm{\delta}_{a}(\bm{y})\right)^{2}\right]-\mathop{\mathbb{E}}_{p(\bm{y}\,|\,\bm{x};\bm{\theta})}\left[\partial_{i}^{2}\bm{h}^{a}(\bm{x})\cdot\bm{\delta}_{a}(\bm{y})\right]^{2}\right)$
$=\frac{1}{N}\cdot\left(\mathop{\mathbb{E}}_{p(\bm{y}\,|\,\bm{x};\bm{\theta})}\left[\partial_{i}^{2}\bm{h}^{a}(\bm{x})\cdot\partial_{i}^{2}\bm{h}^{b}(\bm{x})\cdot\bm{\delta}_{a}(\bm{y})\cdot\bm{\delta}_{b}(\bm{y})\right]-\partial_{i}^{2}\bm{h}^{a}(\bm{x})\cdot\mathop{\mathbb{E}}_{p(\bm{y}\,|\,\bm{x};\bm{\theta})}\left[\bm{\delta}_{a}(\bm{y})\right]^{2}\right)$
$=\frac{1}{N}\cdot\left(\partial_{i}^{2}\bm{h}^{a}(\bm{x})\cdot\partial_{i}^{2}\bm{h}^{b}(\bm{x})\cdot\mathop{\mathbb{E}}_{p(\bm{y}\,|\,\bm{x};\bm{\theta})}\left[\bm{\delta}_{a}(\bm{y})\cdot\bm{\delta}_{b}(\bm{y})\right]\right)$
$=\frac{1}{N}\cdot\partial_{i}^{2}\bm{h}^{a}(\bm{x})\cdot\partial_{i}^{2}\bm{h}^{b}(\bm{x})\cdot\mathcal{I}_{ab}(\bm{h}\,|\,\bm{x}),$
where the second last line follows from the fact that $\bm{\eta}(\bm{x})=\mathop{\mathbb{E}}_{p(\bm{y}\,|\,\bm{x};\bm{\theta})}[\bm{t}(\bm{y})]$ and thus $\mathop{\mathbb{E}}_{p(\bm{y}\,|\,\bm{x};\bm{\theta})}[\bm{\delta}(\bm{y})]=0$. This yields the desired result. ∎
@@
Lemma 3.1 shows that, for the former, ${\mathcal{V}}_{1}(\theta_{i}\,|\,\bm{x})$ only depends on 1st order derivatives; while ${\mathcal{V}}_{2}(\theta_{i}\,|\,\bm{x})$ only depends on the 2nd order derivatives. For the latter, ${\mathcal{V}}_{1}(\theta_{i}\,|\,\bm{x})$ depends on both the 2nd and 4th central moments of $\bm{t}(\bm{y})$; while ${\mathcal{V}}_{2}(\theta_{i}\,|\,\bm{x})$ only depends on the 2nd central moments.
@@
Given $\mathcal{I}_{ab}(\bm{h}\,|\,\bm{x})$ and $\partial_{i}\bm{h}^{a}(\bm{x})$, the computational complexity of all diagonal entries $\mathcal{I}(\theta_{i}\,|\,\bm{x})$ is $\mathcal{O}(T^{2}\dim(\bm{\theta}))$. If $\mathcal{K}^{p}_{abcd}(\bm{t}\,|\,\bm{x})$ and $\partial^{2}_{i}{\bm{h}}^{a}(\bm{x})$ are given, then the computational complexity of the variances in Eqs. 5 and 6 is respectively $\mathcal{O}(T^{4}\dim(\bm{\theta}))$ and $\mathcal{O}(T^{2}\dim(\bm{\theta}))$. Each requires to evaluate a $T\times\dim(\bm{\theta})$ matrix, either $\partial_{i}\bm{h}^{a}(\bm{x})$ or $\partial^{2}_{i}{\bm{h}}^{a}(\bm{x})$ — which can be expensive to calculate for the latter. This is why we need efficient estimators and / or bounds for the tensors on the LHS of Eqs. 4, 5 and 6.
Appendix C Off-Diagonal Variance